OpenStudy (anonymous):

i n33d help pleasssss ;^; The tables below show the values of f(x) and g(x) for different values of x: f(x) = 2(3)x x f(x) -2 0.22 -1 0.67 0 2 1 6 2 18 g(x) = 3x + 1 x g(x) -2 1.11 -1 1.33 0 2 1 4 2 10 Based on the tables, what is the solution to the equation 2(3)x = 3x + 1?

OpenStudy (anonymous):

i guess

OpenStudy (welshfella):

what exactly is 2(3)x mean?

OpenStudy (astrophysics):

Ok so, the solution must have the same x values

OpenStudy (anonymous):

oki

OpenStudy (astrophysics):

In the table what values for the given functions do you have the same values of x in, if that makes sense

OpenStudy (astrophysics):

The (x,y) values must be similar in both of the tables.

OpenStudy (anonymous):

ok so

OpenStudy (astrophysics):

I'm also assuming your function is \[2(3)^x\]

OpenStudy (astrophysics):

\[f(x) = 2(3)^x~~~~g(x) = 3^x+1\] right

OpenStudy (anonymous):

i dunno im sorry

OpenStudy (astrophysics):

Look at your tables of values, plug in your following values for lets say 1, so f(1) = 2(3)^1 = 6 and g(1) = 3^1+1 = 4 so this isn't the one.

OpenStudy (astrophysics):

As the output is not the same, so it's not a solution for your equation...

OpenStudy (anonymous):

oh

OpenStudy (astrophysics):

I feel as if you're not going to do it yourself. So can you plug in the values or something at least so I know I should continue helping you.

OpenStudy (anonymous):

im srry i said it before i need help and i dont understand any of it

OpenStudy (astrophysics):

Can you be more specific, like what don't you understand of what I said? You have to look at your table for both of your equations and basically see which values are similar in both and that will be your solution since if you graph it out, the functions will meet at that point, that's all there is to it.